# Permutation polynomials with a few terms over finite fields

### v2

@@ 4. Permutation polynomials from degree-four rational functions
In this section, we study six classes of permuation pentanomials and two classes of permutation binomials over the finite field $\mathbb{F}_{q^{2}}$ in even characteristic, using degree-four rational functions in Lemma 2.4.
@@
First, we would like to review some background material on linearized polynomials over $\mathbb{F}_{q}$, that would be used to characterise degree-four rational functions $f(X)$ permuting $\mathbb{P}^{1}(\mathbb{F}_{q})$. Polynomials over $\mathbb{F}_{q}$ of the form $L(X)=\sum_{i=0}^{m-1}a_{i}X^{2^{i}}\in\mathbb{F}_{q}[X]$ are often known as additive polynomials or linearized polynomials. Such a special kind of polynomials can induce linear transformations of vector space $\mathbb{F}_{2}^{m}$ over $\mathbb{F}_{2}$. Linearized polynomials are interesting objects especially when they are bijective. More precisely, a linearized polynomial $L(X)$ permutes the finite field $\mathbb{F}_{q}$ if and only if its only root in $\mathbb{F}_{q}$ is zero. The degree-four linearized polynomials $X^{4}+bX^{2}+aX\in\mathbb{F}_{q}[X]$ permutes $\mathbb{F}_{q}$ if and only if $P_{a,b}(X):=X^{3}+bX+a=0$ has no solution in $\mathbb{F}_{q}$. When $(a,b)=(0,0)$ it is clear that $X^{4}$ permutes $\mathbb{F}_{q}$; when $a=0$ and $b\neq 0$, the polynomial $X^{4}+bX^{2}$ is not a permutation of $\mathbb{F}_{q}$ since $P_{a,b}(X)=X(X^{2}+b)$ has a nonzero root in $\mathbb{F}_{q}$; and when $a\neq 0$ and $b=0$, the polynomial $P_{a,b}=X^{4}+aX$ is a permutation of $\mathbb{F}_{q}$ if and only if $a$ is a non-cube element in $\mathbb{F}_{q}$; and lastly, when $a,b\neq 0$, there exists a unique $c\in\mathbb{F}_{q}^{*}$ such that $b=c^{2}$ and by a simple substitution $X=cX$, we can transform $P_{a,b}(X)$ into the form $P_{\alpha}(X)=X^{3}+X+\alpha$. Thus the linearized polynomial $X^{4}+bX^{2}+aX$ is a permutation polynomial if and only if $P_{\alpha}(X)$ is irreducible over $\mathbb{F}_{q}$. The following lemma gives necessary and sufficient conditions for polynomial $P_{\alpha}(X)$ to be irreducible over $\mathbb{F}_{q^{2}}$.
@@
$\nu(X)={\tilde{\beta}}\left(\frac{X+{\tilde{\delta}}^{q}}{X+{\tilde{\delta}}}\right),$
for some ${\tilde{\beta}}\in\mu_{q+1}$ and ${\tilde{\delta}}\in\mathbb{F}_{q^{2}}\backslash\mathbb{F}_{q}$. Now consider the rational function $\nu\circ f\circ\rho:\mu_{q+1}\rightarrow\mu_{q+1}$ which permutes the set $\mu_{q+1}$. The composition is given by
$\begin{split}\nu\circ f\circ\rho~{}(X)&=\nu\circ\left(\frac{(\delta X+\beta\delta^{q})^{4}}{(X+\beta)^{4}}+\frac{(\delta X+\beta\delta^{q})^{2}}{(X+\beta)^{2}}+\frac{\alpha(\delta X+\beta\delta^{q})}{(X+\beta)}\right)\\ &={\tilde{\beta}}\left(\frac{X+{\tilde{\delta}}^{q}}{X+{\tilde{\delta}}}\right)\circ\left(\frac{(\delta X+\beta\delta^{q})^{4}+(\delta X+\beta\delta^{q})^{2}(X+\beta)^{2}+\alpha(\delta X+\beta\delta^{q})(X+\beta)^{3}}{X^{4}+\beta^{4}}\right)\\ &={\tilde{\beta}}\left(\frac{X+{\tilde{\delta}}^{q}}{X+{\tilde{\delta}}}\right)\circ\left(\frac{N_{4}X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+N_{0}}{X^{4}+\beta^{4}}\right)\\ &={\tilde{\beta}}\frac{(N_{4}+{\tilde{\delta}}^{q})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}}^{q})}{(N_{4}+{\tilde{\delta}})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}})},\end{split}$
@@
$\begin{split}h_{1}(X)&=(N_{4}+{\tilde{\delta}})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}}),\\ h_{2}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q}+(N_{4}+{\tilde{\delta}})X^{3}+N_{3}X^{2}+N_{2}X+N_{1},\\ h_{3}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{2q}+N_{1}X^{q}+(N_{4}+{\tilde{\delta}})X^{2}+N_{3}X+N_{2},\\ h_{4}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{3q}+N_{1}X^{2q}+N_{2}X^{q}+(N_{4}+{\tilde{\delta}})X+N_{3},\\ h_{5}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{4q}+N_{1}X^{3q}+N_{2}X^{2q}+N_{3}X^{q}+(N_{4}+{\tilde{\delta}}),\end{split}$
@@
Assume $X^{3}+X+\alpha$ is irreducible over $\mathbb{F}_{q}$. Let $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ such that $\tilde{\delta}\not\in\{\delta^{4}+\delta^{2}+\alpha\delta,\delta^{4q}+\delta^{2q}+\alpha\delta^{q}\}$, $\delta+\delta^{q}+\alpha\neq 0$. Assume $\beta,\tilde{\beta}$ in $\mu_{q+1}$ satisfies ${\tilde{\beta}}\beta^{4}=1$. Then, the pentanomial $f_{1}(X)=X^{4}h_{1}(X^{q-1})$ for $h_{1}(X)=(N_{4}+{\tilde{\delta}})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}})$ is a PP over $\mathbb{F}_{q^{2}}$.
@@
For $q=4$, $f(X)=(b^{2}+b+1)X^{13}+X^{10}+(b^{2}+b+1)X^{7}+(b^{2}+1)X^{4}+(b+1)X$ is a permutation polynomial over the finite field $F_{q^{2}}$, where $b$ is generator of the cyclic group $F_{q^{2}}^{*}$. We have obtained PP by choosing $\beta=1$, $\delta=b^{3}=\tilde{\delta}$ and $\alpha=1$ for the Theorem 4.4. In addition, take $q=2^{8}$, $\beta=1$, $\alpha=b^{q+1}+b^{(q+1)(q-2)}$ and $(\delta,\tilde{\delta})=(b^{321},b^{47351})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{16}+b^{5}+b^{3}+b^{2}+1=0$. Then we obtain from Theorem 4.4 a permutation $f(X)=b^{8722}X^{4q}+b^{48830}X^{3q+1}+b^{53713}X^{2q+2}+b^{48830}X^{q+3}+b^{47311}X^{4}$ over $\mathbb{F}_{q^{2}}$.
@@
Let $h_{5}(X)=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q-3}+N_{1}X^{q-2}+N_{2}X^{q-1}+N_{3}X^{q}+(N_{4}+{\tilde{\delta}}),$ where $N_{0}$, $N_{1},$ $N_{2},$ $N_{3}$ and $N_{4}$ are as given in system (4.1) and $r_{5}=q-3$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $N_{i}$’s, $0\leq i\leq 4$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$, $\tilde{\delta}\not\in\{\delta^{4}+\delta^{2}+\alpha\delta,\delta^{4q}+\delta^{2q}+\alpha\delta^{q}\}$, $\delta+\delta^{q}+\alpha\neq 0$ and $X^{3}+X+\alpha$ is irreducible over $\mathbb{F}_{q}$. Then $f_{5}(X)=X^{r_{5}}h_{5}(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$ when the elements $\beta,\tilde{\beta}\in\mu_{q+1}$ satisfy ${\tilde{\beta}}\beta^{4}+1=0$.
@@
In this subcase, we take the case where $f(X)=X^{4}+\alpha X$, where $\alpha$ is nonzero and non-cube element in $\mathbb{F}_{q}$. Now consider the rational function $\nu\circ f\circ\rho:\mu_{q+1}\rightarrow\mu_{q+1}$ which permutes the set $\mu_{q+1}$ and is given by $\nu\circ f\circ\rho(X)$, which is equal to,
$\tilde{\beta}\frac{X+{\tilde{\delta}}^{q}}{X+\tilde{\delta}}\circ(X^{4}+\alpha X)\circ\frac{\delta X+\beta\delta^{q}}{X+\beta}$
$=$ $\tilde{\beta}\frac{X+{\tilde{\delta}}^{q}}{X+\tilde{\delta}}\circ\left(\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+\alpha\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)\right)$
$=$ $\tilde{\beta}\frac{\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+\alpha\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)+{\tilde{\delta}}^{q}}{\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+\alpha\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)+\tilde{\delta}}$
$=$ $\tilde{\beta}\frac{(\delta^{4}+\alpha\delta+{\tilde{\delta}}^{q})X^{4}+\alpha\beta(\delta+\delta^{q})X^{3}+\alpha\beta^{2}(\delta+\delta^{q})X^{2}+\alpha\beta^{3}(\delta+\delta^{q})X+\beta^{4}(a\delta^{q}+{\tilde{\delta}}^{q}+\delta^{4q})}{(\delta^{4}+\alpha\delta+{\tilde{\delta}})X^{4}+\alpha\beta(\delta+\delta^{q})X^{3}+\alpha\beta^{2}(\delta+\delta^{q})X^{2}+\alpha\beta^{3}(\delta+\delta^{q})X+\beta^{4}(\alpha\delta^{q}+{\tilde{\delta}}+\delta^{4q})}$
@@
(4.2) $\begin{cases}N_{0}&=\beta^{4}(\alpha\delta^{q}+\delta^{4q}),\\ N_{1}&=\alpha\beta^{3}(\delta+\delta^{q}),\\ N_{2}&=\alpha\beta^{2}(\delta+\delta^{q}),\\ N_{3}&=\alpha\beta(\delta+\delta^{q}),\\ N_{4}&=\delta^{4}+\alpha\delta.\end{cases}$
@@
$\begin{split}h_{1}(X)&=(N_{4}+{\tilde{\delta}})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}}),\\ h_{2}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q}+(N_{4}+{\tilde{\delta}})X^{3}+N_{3}X^{2}+N_{2}X+N_{1},\\ h_{3}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{2q}+N_{1}X^{q}+(N_{4}+{\tilde{\delta}})X^{2}+N_{3}X+N_{2},\\ h_{4}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{3q}+N_{1}X^{2q}+N_{2}X^{q}+(N_{4}+{\tilde{\delta}})X+N_{3},\\ h_{5}(X)&=(N_{0}+\beta^{4}{\tilde{\delta}})X^{4q}+N_{1}X^{3q}+N_{2}X^{2q}+N_{3}X^{q}+(N_{4}+{\tilde{\delta}}),\end{split}$
@@
Let $h_{1}(X)=(N_{4}+{\tilde{\delta}})X^{4}+N_{3}X^{3}+N_{2}X^{2}+N_{1}X+(N_{0}+\beta^{4}{\tilde{\delta}})$, where $N_{0}$, $N_{1},$ $N_{2},$ $N_{3}$ and $N_{4}$ are as given in system (4.2) and $r_{1}=4$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $N_{i}$’s, $0\leq i\leq 4$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{4}+\alpha\delta,\delta^{4q}+\alpha\delta^{q}\}$ where $\alpha$ is nonzero and non-cube element in $\mathbb{F}_{q}$. Then $f_{1}(X)=X^{r_{1}}h_{1}(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$ when the elements $\beta,\tilde{\beta}\in\mu_{q+1}$ satisfy ${\tilde{\beta}}\beta^{4}+1=0$.
@@
Let $b$ be a primitive element of $\mathbb{F}_{q^{2}}$ with $q=4$. Take $\beta=b^{3}=\tilde{\beta}$, $\delta=b=\tilde{\delta}$ and $\alpha=b^{2}+b$. From the above Theorem 4.11 we get $f(X)=(b^{2}+1)X^{13}+(b^{3}+b^{2}+b)X^{10}+(b^{3}+1)X^{7}+(b^{3}+b^{2})X^{4}+(b^{3}+b^{2}+1)X$ that is a permutation polynomial over $F_{q^{2}}$. Furthermore, take $q=2^{6}$, $\beta=b^{3087}$, $\alpha=b^{q+1}$, $(\delta,\tilde{\delta})=(b^{3894},b^{3990})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{12}+b^{7}+b^{6}+b^{5}+b^{3}+b+1=0$. Then we obtain from Theorem 4.11 a permutation $f(X)=b^{28}X^{4q}+b^{3477}X^{3q+1}+b^{2469}X^{2q+2}+b^{1461}X^{q+3}+b^{3107}X^{4}$ over $\mathbb{F}_{q^{2}}$.
@@
Let $h_{2}(X)=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q}+(N_{4}+{\tilde{\delta}})X^{3}+N_{3}X^{2}+N_{2}X+N_{1},$ where $N_{0}$, $N_{1},$ $N_{2},$ $N_{3}$ and $N_{4}$ are as given in system (4.2) and $r_{2}=2$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $N_{i}$’s, $0\leq i\leq 4$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{4}+\alpha\delta,\delta^{4q}+\alpha\delta^{q}\}$ where $\alpha$ is nonzero and non-cube element in $\mathbb{F}_{q}$. If ${\tilde{\beta}}\beta^{4}=1$ for some ${\tilde{\beta}}\in\mu_{q+1}$, then $f_{2}(X)=X^{r_{2}}h_{2}(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$.
@@
Let $h_{5}(X)=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q-3}+N_{1}X^{q-2}+N_{2}X^{q-1}+N_{3}X^{q}+(N_{4}+{\tilde{\delta}}),$ where $N_{0}$, $N_{1},$ $N_{2},$ $N_{3}$ and $N_{4}$ are as given in system (4.2) and $r_{5}=q-3$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $N_{i}$’s, $0\leq i\leq 4$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{4}+\alpha\delta,\delta^{4q}+\alpha\delta^{q}\}$ where $\alpha$ is nonzero and non-cube element in $\mathbb{F}_{q}$. Then $f_{5}(X)=X^{r_{5}}h_{5}(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$ when the elements $\beta,\tilde{\beta}\in\mu_{q+1}$ satisfy ${\tilde{\beta}}\beta^{4}+1=0$.
@@
In this subcase, we consider $f(X)=X^{4}$ and give two classes of permutation binomials by putting $\alpha=0$ in the above Subcase 1.2. Using same techniques as above, we summarise the permutation binomials together in the following theorem.
@@
    $f_{5}(X)=X^{q-3}h_{5}(X^{q-1})$, where $h_{5}(X)=(\beta^{4}\delta^{4q}+\beta^{4}{\tilde{\delta}})X^{4q}+(\delta^{4}+{\tilde{\delta}})$, is a permutation binomial over $\mathbb{F}_{q^{2}}$.
@@
Let $q=4$ and $b$ be a primitive element of $\mathbb{F}_{q^{2}}$. By choosing $\beta=b^{3}=\tilde{\beta}$, $\delta=b$ and $\tilde{\delta}=b^{3}$ in the Theorem 4.17, we obtain a permutation polynomial $f(X)=(b^{3}+b^{2})X^{14}+(b^{4}+b^{3})X^{11}$ over $F_{q^{2}}$. Moreover, take $q=2^{8}$, $\beta=b^{31110}$, $(\delta,\tilde{\delta})=(b^{53660},b^{33334})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{16}+b^{5}+b^{3}+b^{2}+1=0$. Then we obtain from Theorem 4.11 a permutation $f(X)=b^{34047}X^{q^{2}-q+2}+b^{53717}X^{3q-1}$ over $\mathbb{F}_{q^{2}}$.
Lemma  2.4 recalled another two cases of degree-four permutation rational functions characterized in [13, Theorem 1.4]:
Case 2:
When $q\leq 8$ and $f(X)$ is equivalent to the rational functions given in Table $1$ [13, Page 19].
Case 3:
When $q$ is odd and $f(X)$ is equivalent to $\dfrac{X^{4}-2\alpha X^{2}-8\beta X+\alpha^{2}}{X^{3}+\alpha X+\beta}$ for some $\alpha,\beta\in\mathbb{F}_{q}$ such that $X^{3}+\alpha X+\beta$ is irreducible in $\mathbb{F}_{q}[X]$.
It appears that handling the above two cases is more computationally cumbersome, especially on the requirement that $\nu\circ f\circ\rho$ should be of the form $X^{r}h(X)^{q-1}$. Better methods are needed to resolve this case.

### v3

@@ 4. Permutation polynomials from degree-four rational functions
In this section, we study six classes of permuation pentanomials and two classes of permutation binomials over the finite field $\mathbb{F}_{q^{2}}$ in even characteristic, using degree-four rational functions given in the first case of Lemma 2.4.
@@
First, we would like to review some background material on linearized polynomials over $\mathbb{F}_{q}$, that would be used to characterise degree-four rational functions $f(X)$ permuting $\mathbb{P}^{1}(\mathbb{F}_{q})$. Polynomials over $\mathbb{F}_{q}$ of the form $L(X)=\sum_{i=0}^{m-1}a_{i}X^{2^{i}}\in\mathbb{F}_{q}[X]$ are often known as additive polynomials or linearized polynomials. Such a special kind of polynomials can induce linear transformations of vector space $\mathbb{F}_{2}^{m}$ over $\mathbb{F}_{2}$. Linearized polynomials are interesting objects especially when they are bijective. More precisely, a linearized polynomial $L(X)$ permutes the finite field $\mathbb{F}_{q}$ if and only if its only root in $\mathbb{F}_{q}$ is zero. The degree-four linearized polynomials $X^{4}+bX^{2}+aX\in\mathbb{F}_{q}[X]$ permutes $\mathbb{F}_{q}$ if and only if $P_{a,b}(X):=X^{3}+bX+a=0$ has no solution in $\mathbb{F}_{q}$. When $(a,b)=(0,0)$ it is clear that $X^{4}$ permutes $\mathbb{F}_{q}$; when $a\neq 0$ and $b=0$, the polynomial $P_{a,b}=X^{4}+aX$ is a permutation of $\mathbb{F}_{q}$ if and only if $a$ is a non-cube element in $\mathbb{F}_{q}$; and lastly, when $b\neq 0$, there exists a unique $c\in\mathbb{F}_{q}^{*}$ such that $b=c^{2}$ and by a simple substitution $X=cX$, we can transform $P_{a,b}(X)$ into the form $P_{\alpha}(X)=X^{3}+X+\alpha$, where $\alpha=\frac{a}{c^{3}}$. Thus the linearized polynomial $X^{4}+bX^{2}+aX$ is a permutation polynomial if and only if $P_{\alpha}(X)$ is irreducible over $\mathbb{F}_{q}$. The following lemma gives necessary and sufficient conditions for polynomial $P_{\alpha}(X)$ to be irreducible over $\mathbb{F}_{q^{2}}$.
@@
[…]
for some ${\tilde{\beta}}\in\mu_{q+1}$ and ${\tilde{\delta}}\in\mathbb{F}_{q^{2}}\backslash\mathbb{F}_{q}$. Now consider the rational function $\nu\circ f\circ\rho:\mu_{q+1}\rightarrow\mu_{q+1}$ which permutes the set $\mu_{q+1}$. The composition $\nu\circ f\circ\rho~(X)$ is given by
[…]
@@
[…]
@@
[…]
@@
For $q=4$, $f(X)=(b^{2}+b+1)X^{13}+X^{10}+(b^{2}+b+1)X^{7}+(b^{2}+1)X^{4}+(b+1)X$ is a permutation polynomial over the finite field $\mathbb{F}_{q^{2}}$, where $b$ is generator of the cyclic group $\mathbb{F}_{q^{2}}^{*}$. We have obtained $P$ by choosing $\beta=1$, $\delta=b^{3}=\tilde{\delta}$ and $\alpha=1$ for the Theorem 4.4. In addition, take $q=2^{8}$, $\beta=1$, $\alpha=b^{q+1}+b^{(q+1)(q-2)}$ and $(\delta,\tilde{\delta})=(b^{321},b^{47351})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{16}+b^{5}+b^{3}+b^{2}+1=0$. Then we obtain from Theorem 4.4 a permutation $f(X)=b^{8722}X^{4q}+b^{48830}X^{3q+1}+b^{53713}X^{2q+2}+b^{48830}X^{q+3}+b^{47311}X^{4}$ over $\mathbb{F}_{q^{2}}$.
@@
[…]
@@
In this subcase, we take the case where $f(X)=X^{4}+aX$, where $a$ is nonzero and non-cube element in $\mathbb{F}_{q}$. Now consider the rational function $\nu\circ f\circ\rho:\mu_{q+1}\rightarrow\mu_{q+1}$ which permutes the set $\mu_{q+1}$ and is given by $\nu\circ f\circ\rho(X)$, which is equal to,
$\tilde{\beta}\frac{X+{\tilde{\delta}}^{q}}{X+\tilde{\delta}}\circ(X^{4}+aX)\circ\frac{\delta X+\beta\delta^{q}}{X+\beta}$
$=$ $\tilde{\beta}\frac{X+{\tilde{\delta}}^{q}}{X+\tilde{\delta}}\circ\left(\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+a\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)\right)$
$=$ $\tilde{\beta}\frac{\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+a\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)+{\tilde{\delta}}^{q}}{\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)^{4}+a\left(\frac{\delta X+\beta\delta^{q}}{X+\beta}\right)+\tilde{\delta}}$
$=$ $\tilde{\beta}\frac{(\delta^{4}+a\delta+{\tilde{\delta}}^{q})X^{4}+a(\delta+\delta^{q})\left(\beta X^{3}+\beta^{2}X^{2}+\beta^{3}X\right)+\beta^{4}(a\delta^{q}+{\tilde{\delta}}^{q}+\delta^{4q})}{(\delta^{4}+a\delta+{\tilde{\delta}})X^{4}+a(\delta+\delta^{q})\left(\beta X^{3}+\beta^{2}X^{2}+\beta^{3}X\right)+\beta^{4}(a\delta^{q}+{\tilde{\delta}}+\delta^{4q})}$
@@
(4.2) $\begin{cases}N_{0}&=\beta^{4}(a\delta^{q}+\delta^{4q}),\\ N_{1}&=a\beta^{3}(\delta+\delta^{q}),\\ N_{2}&=a\beta^{2}(\delta+\delta^{q}),\\ N_{3}&=a\beta(\delta+\delta^{q}),\\ N_{4}&=\delta^{4}+a\delta.\end{cases}$
@@
[…]
@@
[…]
@@
Let $b$ be a primitive element of $\mathbb{F}_{q^{2}}$ with $q=4$. Take $\beta=b^{3}=\tilde{\beta}$, $\delta=b=\tilde{\delta}$ and $a=b^{2}+b$. From the above Theorem 4.11 we get $f(X)=(b^{2}+1)X^{13}+(b^{3}+b^{2}+b)X^{10}+(b^{3}+1)X^{7}+(b^{3}+b^{2})X^{4}+(b^{3}+b^{2}+1)X$ that is a permutation polynomial over $\mathbb{F}_{q^{2}}$. Furthermore, take $q=2^{6}$, $\beta=b^{3087}$, $a=b^{q+1}$, $(\delta,\tilde{\delta})=(b^{3894},b^{3990})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{12}+b^{7}+b^{6}+b^{5}+b^{3}+b+1=0$. Then we obtain from Theorem 4.11 a permutation $f(X)=b^{28}X^{4q}+b^{3477}X^{3q+1}+b^{2469}X^{2q+2}+b^{1461}X^{q+3}+b^{3107}X^{4}$ over $\mathbb{F}_{q^{2}}$.
@@
Let $h_{2}(X)=(N_{0}+\beta^{4}{\tilde{\delta}})X^{q}+(N_{4}+{\tilde{\delta}})X^{3}+N_{3}X^{2}+N_{2}X+N_{1},$ where $N_{0}$, $N_{1},$ $N_{2},$ $N_{3}$ and $N_{4}$ are as given in system (4.2) and $r_{2}=2$. Moreover, $\delta$ and $\tilde{\delta}$ in the expressions of $N_{i}$’s, $0\leq i\leq 4$ are such that $\delta,\tilde{\delta}\in\mathbb{F}_{q^{2}}\setminus\mathbb{F}_{q}$ and $\tilde{\delta}\not\in\{\delta^{4}+a\delta,\delta^{4q}+a\delta^{q}\}$ where $a$ is nonzero and non-cube element in $\mathbb{F}_{q}$. If ${\tilde{\beta}}\beta^{4}=1$ for some ${\tilde{\beta}}\in\mu_{q+1}$, then $f_{2}(X)=X^{r_{2}}h_{2}(X^{q-1})$ is a PP over $\mathbb{F}_{q^{2}}$.
@@
[…]
@@
In this subcase, we consider $f(X)=X^{4}$ and give two classes of permutation binomials by putting $a=0$ in the above Subcase 1.2. Using same techniques as above, we summarise the permutation binomials together in the following theorem.
@@
[…]
@@
Let $q=4$ and $b$ be a primitive element of $\mathbb{F}_{q^{2}}$. By choosing $\beta=b^{3}=\tilde{\beta}$, $\delta=b$ and $\tilde{\delta}=b^{3}$ in the Theorem 4.17, we obtain a permutation polynomial $f(X)=(b^{3}+b^{2})X^{14}+(b^{4}+b^{3})X^{11}$ over $\mathbb{F}_{q^{2}}$. Moreover, take $q=2^{8}$, $\beta=b^{31110}$, $(\delta,\tilde{\delta})=(b^{53660},b^{33334})$, where $b$ is a primitive element of $\mathbb{F}_{q^{2}}$ satisfying $b^{16}+b^{5}+b^{3}+b^{2}+1=0$. Then we obtain from Theorem 4.11 a permutation $f(X)=b^{34047}X^{q^{2}-q+2}+b^{53717}X^{3q-1}$ over $\mathbb{F}_{q^{2}}$.
Note that Lemma 2.4 consists of two additional cases of degree-four permutation rational functions as characterized in [13, Theorem 1.4]. However, these cases are more computationally demanding, primarily due to the requirement that the composition $\nu\circ f\circ\rho$ must take the form $X^{r}h(X)^{q-1}$ for some $h(X)\in\mathbb{F}_{q^{2}}[X]$. The development of more effective methods to handle these cases remains an open problem and a promising direction for further research.